OpenStudy (anonymous):

How do you solve a quadratic equation using the quadratic formula? Could you demonstrate the process?

OpenStudy (anonymous):

Ok, quadratic equations have the following form: ax^2+bx+c=0, right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, first you need to calculate b^2-4ac. Let's call it D. This number tells you how many solutions exist. if D>0 there are two solutions, if D=0 there is only one solution and if D<0 there are none.

OpenStudy (phi):

If you need more info, watch http://www.khanacademy.org/math/algebra/quadtratics/v/quadratic-formula-1

OpenStudy (anonymous):

Now you can use the quadratic formula to get the solutions: \[x_1=\frac{-b-\sqrt{D}}{2a}\] \[x_2=\frac{-b+\sqrt{D}}{2a}\]

OpenStudy (anonymous):

Note that this doesn't exist if D<0, because you can't take the square root of a negative number. Also note that x1=x2 if D=0, so there's indeed only one solution.

OpenStudy (anonymous):

Ok now I'm beginning to understand this a lot better. Thank

hero (hero):

popjump, give us of one of the problems you are trying to solve using the quadratic formula

OpenStudy (anonymous):

Let me finish this staff meeting and I will one of the problems.

Directrix (directrix):

Here is an example. There are more at the link. http://www.purplemath.com/modules/quadform.htm